how do you write 6(2/3) in simplest form
Found 2 solutions by Alan3354, tinbar:
Answer by Alan3354(69443)

6*(2/3) = 12/3 = 4

Answer by tinbar(133)

6*2 = 12, 12/3 = 4
OR 6/3 = 2, 2*2 = 4
Either way, we get the simplest answer as 4